OpenStudy (anonymous):

solve for lim √(x-2)² divide x-2...using left and right hand limit

OpenStudy (anonymous):

Can be rewritten as \[\lim _{x \rightarrow2+}\left| x-2 \right|/(x-2) = 1.\] \[\lim _{x \rightarrow 2-}\left| x - 2 \right|/(x - 2) = -1.\]